# xOffense: An AI-driven autonomous penetration testing framework with offensive knowledge-enhanced LLMs and multi agent systems

### v1

@@ 2.6 Positioning of our work
Relative to single-agent PentestGPT [7] and post-breach AutoAttacker [18], we retain a multi-agent/PTG discipline akin to PentestAgent/VulnBot [8, 9] but differ in three ways: (i) prioritizing mid-scale, open backbones for cost-effective, on-prem deployment; (ii) employing grey-box phase prompting to maintain phase continuity while limiting drift; and (iii) aligning evaluation with open substrates (AutoPenBench, AI-Pentest-Benchmark, and where applicable CVE-Bench) under fixed budgets and subtask breakdowns [10, 11, 26, 27].
@@
Outcomes hinge on (i) grounding quality (RAG/summaries/validators) and (ii) orchestration discipline (roles/PTG/reflection). Single-agent pipelines set baselines; role-structured multi-agent systems improve reliability; reflective and IP-to-shell variants push autonomy at kill-chain ends; one-day/zero-day studies quantify limits of discovery vs. exploitation [22, 23]. Our approach emphasizes cost-effective, reproducible deployment with mid-scale open models, PTG structure, and grey-box prompts under open, reproducible protocols [10, 11, 26, 27].
@@ 3 Background
3.1 Automated Penetration Testing
Penetration testing aims to evaluate the security of a target system by simulating adversarial behavior across phases such as reconnaissance, vulnerability enumeration, exploitation, and privilege escalation. Manual execution is effective but limited by human resources and scalability. Automated Penetration Testing (APT) addresses these limitations by orchestrating these phases through intelligent agents and machine learning models.
@@ 3.2 Multi-Agent AI Systems
Multi-Agent Systems (MAS) provide a natural architecture for automated penetration testing by assigning each pentest phase to a specialized agent. For example:
@@ 3.3 LLM-based Offensive Agents
Within MAS, LLMs provide reasoning, contextual understanding, and code synthesis capabilities that align well with penetration testing workflows. The central problem is, given an attack context $C$ (system description, logs, CVEs), generate actionable steps or payloads $A$ that maximize the likelihood of successful exploitation:
@@
Although proprietary LLMs (e.g., GPT-4, Claude) offer strong performance, they introduce limitations in cost, reproducibility, and security control. Therefore, we adopt open-source models such as Qwen3, which support local deployment, fine-tuning, and quantization (AWQ/INT4), making them suitable for offensive research environments where sensitive data cannot leave the infrastructure.
@@ 3.4 Fine-tuning Methods for Offensive LLMs
Adapting LLMs to offensive security tasks requires specialization beyond general pre-trained knowledge, and several parameter-efficient fine-tuning approaches have been explored, including Prefix-Tuning, Adapter-based methods, Low-Rank Adaptation (LoRA) and its extension QLoRA.
Prefix-Tuning or P-Tuning v2 appends task-specific continuous vectors or tokens to the input prompt. Its advantages lie in simplicity and a very low memory footprint, since only the prefix embeddings are optimized. However, this method often struggles to capture deeper structural knowledge, and its effectiveness diminishes when reasoning requires multi-step tool interaction or long-context planning, both of which are essential in penetration testing.
@@
Considering the need to adapt a 32B-parameter model such as Qwen3 under realistic hardware constraints, this work adopts LoRA fine-tuning. This approach makes it possible to embed offensive knowledge—covering vulnerability patterns, exploit reasoning, and payload generation—efficiently while preserving the base model’s general-purpose capabilities. Detailed dataset construction and the training procedure are described in Section 4.
@@ 4.1 Overview of the proposed framework
xOffense is an innovative, lightweight framework for autonomous penetration testing, engineered to replicate the collaborative dynamics of human security teams while operating within resource-constrained environments. By harnessing compact Large Language Models (LLMs) with approximately 32 billion parameters, xOffense eliminates dependency on commercial APIs, enabling deployment on standard hardware. The framework decomposes the intricate process of penetration testing into three meticulously designed phases—reconnaissance, scanning, and exploitation - coordinated through a sophisticated multi-agent architecture. Comprising five core components - Task Orchestrator, Knowledge Repository, Command Synthesizer, Action Executor, and Information Aggregator - xOffense ensures seamless task progression, robust information management, and precise execution. This section elucidates the system’s architecture, role delineation, task coordination, inter-agent communication, and execution mechanisms, underscoring its efficacy in addressing cybersecurity challenges with minimal computational overhead.
@@
The operational workflow of xOffense, illustrated in Figure 1, initiates when a user submits a penetration testing objective, such as “Identify vulnerabilities on IP 192.168.X.X and retrieve root-level flags.” This task description serves as the Initial Context, which is passed to the Task Orchestrator for comprehensive plan generation. The orchestrator constructs a Task Coordination Graph (TCG), decomposing the penetration objective into a structured sequence of tasks with clearly defined dependencies. To enhance contextual accuracy, it queries the Knowledge Repository—a vector-based database—via a Retrieval - Augmented Generation (RAG) mechanism, which is taken from Langchain-Chatchat [30], retrieving relevant penetration knowledge based on inputs such as the initial task description, the current task’s instruction, or recent task results.
@@
Each task within the TCG is processed through an iterative loop involving Command Synthesis, Execution, Feedback Analysis, and Dynamic Plan Update. The Command Synthesizer, fine-tuned using lightweight LoRA techniques, translates task directives into precise, tool-specific commands, which are executed by the Action Executor utilizing a MemAgent - enhanced context management system to handle verbose outputs effectively. Post-execution, task outcomes are evaluated and relayed back to the orchestrator, which marks tasks as completed or triggers reflection and re-planning in case of failures. Upon completing all tasks within a phase (e.g., Reconnaissance), the Information Aggregator consolidates outputs into concise directives for the subsequent phase (e.g., Scanning), ensuring contextual coherence and minimizing token overhead. This orchestration - execution-feedback loop is consistently applied across all three phases, with each phase iteratively refining the attack path based on task outcomes and environmental feedback. The overall workflow, including execution and re-planning, is later formalized in Algorithm 1. The workflow terminates upon successfully achieving the defined success criteria, such as privilege escalation or flag retrieval.
@@ 4.2 Role Specialization
To navigate the complexity of penetration testing, xOffense employs a role specialization strategy, mitigating the risk of information overload and ensuring contextual coherence across phases. By assigning agents to distinct roles, the framework optimizes resource utilization and maintains precision in task execution, addressing the challenge of dynamic reasoning across testing stages.
@@ 4.3 Task Coordination and Reflection
The Task Coordination Graph (TCG) and its integrated Check and Reflection Mechanism form the cornerstone of xOffense’s penetration path planning, enabling systematic task execution and adaptive plan refinement. These components address the challenges of limited context windows and inadequate error handling, ensuring robust and dynamic testing workflows. Their operational logic is illustrated in Algorithm 1, Algorithm 2, and Algorithm 3.
@@ 4.3.1 Task Coordination Graph
    Directive: A clear instruction, e.g., ”enumerate services on port 80 of 192.168.X.X.”
@@
    Operation Type: Specifies whether the task involves automated shell commands (e.g., nmap) or manual intervention.
@@
The Task Orchestrator generates the TCG in a JSON - compliant format, dynamically updating it based on execution outcomes. For instance, a task designed to perform user authentication by initiating an SSH connection attempt—targeting a specific service endpoint associated with remote shell access protocols—on port 22 of 192.168.X.X depends on the successful completion of a preceding port scanning task. Subsequent tasks, such as performing an exhaustive enumeration of writable directories for privilege escalation through misconfigured permissions or publicly writable paths — (find / -writable 2>/dev/null) or listing running processes (ps aux), are contingent on this authentication.
@@
Figure 2: Task Coordination Graph (TCG) illustrating task dependencies and execution status. Completed tasks are shown in dark, the current task in orange, and pending tasks in light blue.
@@ 4.3.2 Check and Reflection Mechanism
The generation of erroneous commands and the lack of effective error-handling mechanisms pose significant challenges to LLM-based penetration testing. xOffense addresses these issues through a Check and Reflection Mechanism integrated into the Task Orchestrator, enabling continuous self-assessment and plan optimization. The full workflow is detailed in Algorithm 1.
During the Task Session, the Action Executor evaluates task outcomes and updates the Task Coordination Graph (TCG) with success or failure statuses. The Planning Session then reflects on these outcomes, revising task directives and updating the TCG accordingly. Successful tasks are retained, while failed tasks trigger a reanalysis process, wherein the LLM regenerates commands with corrected parameters or alternative strategies. The updated plan is merged with previously completed tasks to preserve execution continuity, as demonstrated in Algorithm 2 and Algorithm 3.
@@
1:TCG (Task Coordination Graph), Knowledge Repository KRKR
@@
The Knowledge Repository supports this mechanism indirectly by assisting the Task Orchestrator during plan updates. It stores embeddings of previously successful tasks and curated penetration testing knowledge (e.g., exploitation techniques, privilege escalation methods, and tool usage tutorials from sources such as HackTricks [38] and HackingArticles [39]). When re-planning, the Task Orchestrator queries this repository to retrieve the top-$k$ most relevant past cases using vector similarity search. Retrieved results are re-ranked and integrated into the revised TCG, ensuring that updates benefit from prior successes. This integration enhances resilience against hallucinated commands, improves error recovery, and maintains efficiency across iterative penetration testing phases.
@@ 4.4 Inter-Agent Communication Mechanism
For example, reconnaissance outputs—such as open ports (e.g., 22, 80, 443), service versions, and system fingerprints—are synthesized into a compact directive for the scanning phase, enabling targeted vulnerability detection with tools like Nikto or sqlmap. Similarly, scanning outputs, such as a SQL injection vulnerability identified by sqlmap or a misconfiguration detected by Nuclei, are summarized to guide the exploitation phase in prioritizing relevant exploits.
The Information Aggregator maintains a persistent shell state log, tracking access levels (e.g., a low-privileged user account gained via SSH) and system context (e.g., operating system type). This log ensures continuity across phases, mitigating the risk of context loss and enabling dynamic integration of findings. By filtering outputs from preceding phases to focus on critical insights, the mechanism minimizes computational overhead, ensuring efficient operation on a 32B-parameter LLM. This streamlined communication fosters a cohesive testing process, addressing the challenge of synthesizing information across multiple stages.
@@ 4.5 Generative Behavior and Execution
xOffense supports three operational modes—automatic, semi-automatic, and manual—with the automatic mode enabling fully autonomous testing. The Command Synthesizer transforms TCG directives into precise, tool-specific instructions tailored to the target system and phase, addressing the challenge of accurate command generation.
@@
The Action Executor runs these commands via a Python Paramiko-based interactive shell on a Kali Linux environment, simulating human interactions with high fidelity. This component seamlessly processes tool-specific instructions generated by the Command Synthesizer, enabling robust interaction with the target system through simulated keyboard operations. The Action Executor is optimized for the Qwen3-32B model, which, despite its constrained 16,384-token context window, integrates the innovative MemAgent [40] framework to handle extended contexts effectively. Drawing on MemAgent’s segment-based processing and reinforcement learning (RL)-optimized memory mechanism, as described in the referenced study, the Action Executor processes arbitrarily long outputs by iteratively reading command results in chunks and updating a fixed-length memory. This approach ensures linear computational complexity, allowing xOffense to manage verbose tool outputs without performance degradation, even beyond the Qwen3-32B’s native context limit. To address the challenge of excessive or redundant output, a sophisticated filtering mechanism employs the MemAgent-enhanced LLM to extract critical information when results exceed 8,000 characters, preserving only actionable insights for analysis. These insights are relayed to the Task Orchestrator for further processing, ensuring contextual coherence and minimizing computational overhead. By incorporating MemAgent’s ability to selectively retain relevant data while discarding distractors, the Action Executor enhances the system’s efficiency and scalability, enabling robust penetration testing in resource-constrained environments and contributing to xOffense’s capability to handle complex, long-context workflows with precision.
@@ 5 Implementation, Benchmark Dataset and Metrics
To thoroughly assess the effectiveness and practicality of xOffense, we designed a comprehensive experimental setup that evaluates not only task completion rates but also the scalability, adaptability, and efficiency of the system under realistic penetration testing conditions. This section details the experimental settings, models, fine-tuning methodologies, benchmark datasets, evaluation metrics and presents an in-depth analysis of the empirical results.
@@ 5.1 Experimental Setup
5.1.1 Environment
Attacker Environment
All penetration testing experiments were executed from a dedicated attacker machine configured as a VMware virtual workstation running Kali Linux 2025 [41]. Kali was chosen for its comprehensive penetration testing toolkit and compatibility with industry-standard workflows. This virtualized attacker host was provisioned with 8 vCPUs, 16 GB RAM, and 120 GB storage, ensuring stable execution of both offensive tools and the xOffense multi-agent framework within a single environment.
Victim Environment
Two types of target environments were deployed corresponding to the benchmark datasets:

### v2

@@ 2.6 Positioning of our work
Relative to single-agent PentestGPT [5] and post-breach AutoAttacker [44], we retain a multi-agent/PTG discipline akin to PentestAgent/VulnBot [34, 17] but differ in three ways: (i) prioritizing mid-scale, open backbones for cost-effective, on-prem deployment; (ii) employing grey-box phase prompting to maintain phase continuity while effectively limiting drift; and (iii) aligning evaluation with open substrates (AutoPenBench, AI-Pentest-Benchmark, and where applicable CVE-Bench) under fixed budgets and sub-task breakdowns [9, 16, 47, 15].
@@
Outcomes hinge on (i) grounding quality (RAG/summaries/validators) and (ii) orchestration discipline (roles/PTG/reflection). Single-agent pipelines set baselines; role-structured multi-agent systems consistently improve reliability; reflective and IP-to-shell variants further push autonomy at kill-chain ends; one-day/zero-day studies quantify limits of discovery vs. exploitation [7, 48]. Our approach emphasizes cost-effective, reproducible deployment with mid-scale open models, PTG structure, and grey-box prompts under open, reproducible protocols [9, 16, 47, 15].
@@ 3 Background
3.1 Automated Pentest
Pentest aims to evaluate the security of a target system by simulating adversarial behavior across phases such as reconnaissance, vulnerability enumeration, exploitation, and privilege escalation. Manual execution is effective but limited by human resources and scalability. Automated Pentest (APT) addresses these limitations by orchestrating these phases through intelligent agents and machine learning models.
@@ 3.2 Multi-Agent AI Systems
Multi-Agent Systems (MAS) provide a natural architecture for automated pentest by assigning each pentest phase to a specialized agent. For example:
@@ 3.3 LLM-based Offensive Agents
Within MAS, LLMs provide reasoning, contextual understanding, and code synthesis capabilities that align well with pentest workflows. The central problem is, given an attack context $C$ (system description, logs, CVEs), generate actionable steps or payloads $A$ that maximize the likelihood of successful exploitation:
@@
Although proprietary LLMs, such as GPT-4 and Claude, offer strong performance, they introduce limitations in cost, reproducibility, and security control. Therefore, we adopt open-source models such as Qwen3, which support local deployment, fine-tuning, and quantization (AWQ/INT4), making them suitable for offensive research environments where sensitive data cannot leave the infrastructure.
@@ 3.4 Fine-tuning Methods for Offensive LLMs
Adapting LLMs to offensive security tasks requires task-specific specialization beyond general pre-trained knowledge, and several parameter-efficient fine-tuning approaches have been explored, including Prefix-Tuning, Adapter-based methods, Low-Rank Adaptation (LoRA) and its extension QLoRA.
Prefix-Tuning or P-Tuning v2 appends task-specific continuous vectors or tokens to the input prompt. Its advantages lie in simplicity and a very low memory footprint, since only the prefix embeddings are optimized. However, this method often struggles to capture deeper structural knowledge, and its effectiveness diminishes when reasoning requires multi-step tool interaction or long-context planning, both of which are essential in pentest.
@@
Considering the need to adapt a 32B-parameter model such as Qwen3 under realistic hardware constraints, this work adopts LoRA fine-tuning. This approach makes it possible to embed offensive knowledge, including vulnerability patterns, exploit reasoning, and payload generation, efficiently while preserving the base model’s general-purpose capabilities. Detailed dataset construction and the training procedure are described in Section 4.
@@ 4.1 Overview of the proposed framework
xOffense is an innovative, lightweight framework for autonomous pentest, engineered to replicate the collaborative dynamics of human security teams while operating within resource-constrained environments. By harnessing compact LLMs with approximately 32 billion parameters, xOffense eliminates dependency on commercial APIs, enabling deployment on standard hardware. The framework decomposes the intricate process of pentest into three meticulously designed phases reconnaissance, scanning, and exploitation coordinated through a sophisticated multi-agent architecture. Comprising five core components Task Orchestrator, Knowledge Repository, Command Synthesizer, Action Executor, and Information Aggregator. xOffense ensures seamless task progression, robust information management, and precise execution. This section elucidates the system’s architecture, role delineation, task coordination, inter-agent communication, and execution mechanisms, underscoring its efficacy in addressing cybersecurity challenges with minimal computational overhead.
@@
The operational workflow of xOffense, illustrated in Figure 1, initiates when a user submits a pentest objective, such as “Identify vulnerabilities on IP 192.168.X.X and retrieve root-level flags.” This task description serves as the Initial Context, which is passed to the Task Orchestrator for comprehensive plan generation. The orchestrator constructs a Task Coordination Graph (TCG), decomposing the penetration objective into a structured sequence of tasks with clearly defined dependencies. To enhance contextual accuracy, it queries the Knowledge Repository a vector-based database via a Retrieval-Augmented Generation (RAG) mechanism, which is taken from Langchain-Chatchat [19], retrieving relevant penetration knowledge based on inputs such as the initial task description, the current task’s instruction, or recent task results.
@@
Each task within the TCG is processed through an iterative and adaptive loop involving Command Synthesis, Execution, Feedback Analysis, and Dynamic Plan Update. The Command Synthesizer, fine-tuned using lightweight LoRA techniques, translates task directives into precise, tool-specific commands, which are executed by the Action Executor utilizing a MemAgent-enhanced context management system to handle verbose outputs effectively. Post-execution, task outcomes are evaluated and relayed back to the orchestrator, which marks tasks as completed or triggers reflection and re-planning in case of failures. Upon completing all tasks within a phase, such as Reconnaissance, the Information Aggregator consolidates outputs into concise directives for the subsequent phase, such as Scanning, ensuring contextual coherence and minimizing token overhead. This orchestration-execution-feedback loop is consistently applied across all three phases, with each phase iteratively refining the attack path based on task outcomes and environmental feedback. The overall workflow, including execution and re-planning, is later formalized in Algorithm 1. The workflow terminates upon successfully achieving the defined success criteria, such as privilege escalation or flag retrieval.
@@ 4.2 Role Specialization
To navigate the complexity of pentest, xOffense employs a role specialization strategy, mitigating the risk of information overload and ensuring contextual coherence across phases. By assigning agents to distinct roles, the framework optimizes resource utilization and maintains precision in task execution, addressing the challenge of dynamic reasoning across testing stages.
@@ 4.3 Task Coordination and Reflection
The TCG and its integrated Check and Reflection Mechanism form the cornerstone of xOffense’s penetration path planning, enabling systematic task execution and adaptive plan refinement. These components address the challenges of limited context windows and inadequate error handling, ensuring robust and dynamic testing workflows. Their operational logic is illustrated in Algorithm 1, Algorithm 2, and Algorithm 3.
@@ 4.3.1 Task Coordination Graph
    Directive: A clear instruction, such as ”enumerate services on port 80 of 192.168.X.X.”
@@
    Operation Type: Specifies whether the task involves automated shell commands, such as nmap, or manual intervention.
@@
The Task Orchestrator generates the TCG in a JSON, which is compliant format, dynamically updating it based on execution outcomes. For instance, a task designed to perform user authentication by initiating an SSH connection attempt, targeting a specific service endpoint associated with remote shell access protocols, on port 22 of 192.168.X.X depends on the successful completion of a preceding port scanning task. Subsequent tasks, such as performing an exhaustive enumeration of writable directories for privilege escalation through misconfigured permissions or publicly writable paths (find / -writable 2>/dev/null) or listing running processes (ps aux), are contingent on this authentication.
@@
Figure 2: TCG illustrating task dependencies and execution status. Completed tasks are shown in dark, the current task in orange, and pending tasks in light blue.
@@ 4.3.2 Check and Reflection Mechanism
The generation of erroneous commands and the lack of effective error-handling mechanisms pose significant challenges to LLM-based pentest. xOffense addresses these issues through a Check and Reflection Mechanism integrated into the Task Orchestrator, enabling continuous self-assessment and plan optimization. The full workflow is detailed in Algorithm 1.
During the Task Session, the Action Executor evaluates task outcomes and updates the TCG with success or failure statuses. The Planning Session then reflects on these outcomes, revising task directives and updating the TCG accordingly. Successful tasks are retained, while failed tasks trigger a reanalysis process, wherein the LLM regenerates commands with corrected parameters or alternative strategies. The updated plan is merged with previously completed tasks to preserve execution continuity, as demonstrated in Algorithm 2 and Algorithm 3.
@@
1:TCG, Knowledge Repository KRKR
@@
Plan Update and Merge Algorithms.
The Knowledge Repository supports this mechanism indirectly by assisting the Task Orchestrator during plan updates. It stores embeddings of previously successful tasks and curated pentest knowledge, including exploitation techniques, privilege escalation methods, and tool usage tutorials from sources such as HackTricks [1] and HackingArticles [29]. When re-planning, the Task Orchestrator queries this repository to retrieve the top-$k$ most relevant past cases using vector similarity search. Retrieved results are re-ranked and integrated into the revised TCG, ensuring that updates benefit from prior successes. This integration enhances resilience against hallucinated commands, improves error recovery, and maintains efficiency across iterative pentest phases.
@@ 4.4 Inter-Agent Communication Mechanism
For example, reconnaissance outputs, such as open ports (such as 22, 80, and 443), service versions, and system fingerprints, are synthesized into a compact directive for the scanning phase, enabling targeted vulnerability detection with tools like Nikto or sqlmap. Similarly, scanning outputs, such as a SQL injection vulnerability identified by sqlmap or a misconfiguration detected by Nuclei, are summarized to guide the exploitation phase in prioritizing relevant exploits.
The Information Aggregator maintains a persistent shell state log, tracking access levels, such as a low-privileged user account gained via SSH, and system context, such as operating system type. This log ensures continuity across phases, mitigating the risk of context loss and enabling dynamic integration of findings. By filtering outputs from preceding phases to focus on critical insights, the mechanism minimizes computational overhead, ensuring efficient operation on a 32B-parameter LLM. This streamlined communication fosters a cohesive testing process, addressing the challenge of synthesizing information across multiple stages.
@@ 4.5 Generative Behavior and Execution
xOffense supports three operational modes, specifically automatic, semi-automatic, and manual, with the automatic mode enabling fully autonomous testing. The Command Synthesizer transforms TCG directives into precise, tool-specific instructions tailored to the target system and phase, addressing the challenge of accurate command generation.
@@
The Action Executor runs these commands via a Python Paramiko-based interactive shell on a Kali Linux environment, simulating human interactions with high fidelity. This component seamlessly processes tool-specific instructions generated by the Command Synthesizer, enabling robust interaction with the target system through simulated keyboard operations. The Action Executor is optimized for the Qwen3-32B model, which, despite its constrained 16,384-token context window, integrates the innovative MemAgent [46] framework to handle extended contexts effectively. Drawing on MemAgent’s segment-based processing and reinforcement learning (RL)-optimized memory mechanism, as described in the referenced study, the Action Executor processes arbitrarily long outputs by iteratively reading command results in chunks and updating a fixed-length memory. This approach ensures linear computational complexity, allowing xOffense to manage verbose tool outputs without performance degradation, even beyond the Qwen3-32B’s native context limit. To address the challenge of excessive or redundant output, a sophisticated filtering mechanism employs the MemAgent-enhanced LLM to extract critical information when results exceed 8,000 characters, preserving only actionable insights for analysis. These insights are relayed to the Task Orchestrator for further processing, ensuring contextual coherence and minimizing computational overhead. By incorporating MemAgent’s ability to selectively retain relevant data while discarding distractors, the Action Executor enhances the system’s efficiency and scalability, enabling robust pentest in resource-constrained environments and contributing to xOffense’s capability to handle complex, long-context workflows with precision.
@@ 5 Implementation, Benchmark Dataset and Metrics
To thoroughly assess the effectiveness and practicality of xOffense, we designed a comprehensive experimental setup that evaluates not only task completion rates but also the scalability, adaptability, and efficiency of the system under realistic pentest conditions. This section details the experimental settings, models, fine-tuning method, benchmark datasets, evaluation metrics and presents an in-depth analysis of the empirical results.
@@ 5.1 Experimental Setup
5.1.1 Attacker Environment
All pentest experiments are executed from a dedicated attacker machine configured as a VMware virtual workstation running Kali Linux 2025 [25]. Kali is chosen for its comprehensive pentest toolkit and compatibility with industry-standard workflows. This virtualized attacker host is provisioned with 8 vCPUs, 16 GB RAM, and 120 GB storage, ensuring stable execution of both offensive tools and the xOffense multi-agent framework within a single environment.
5.1.2 Victim Environment
Two types of target environments are deployed corresponding to the benchmark datasets: